OpenStudy (anonymous):

What is the distance between the points (1,4) and (4,8). can someone help me?

OpenStudy (anonymous):

square root of (x2 -x1)2 + (y2 - Y1) 2

OpenStudy (anonymous):

ok so you would get Square root of (4-1)^2 + (8-4)^2

OpenStudy (anonymous):

do you think you can do the rest or do you need the answer?

OpenStudy (anonymous):

hang on. kinda still sleeping.

OpenStudy (anonymous):

ok....I'll wait. :)

OpenStudy (anonymous):

so 9+14

OpenStudy (anonymous):

then 23.

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

thats not what i got...do you have multiple choice?

OpenStudy (anonymous):

nope its true or false lol

OpenStudy (anonymous):

it says its the distance 13

OpenStudy (anonymous):

ok...well i got 5 as an answer

OpenStudy (anonymous):

How the heck. I need to figure this out so I can do it on my EOC later on.

OpenStudy (anonymous):

i might be wrong though...hang on

pooja195 (pooja195):

Start with the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] plug in the info \[\huge d=\sqrt{(4-1)^2+(8-4)^2}\]

OpenStudy (anonymous):